# arXiv:2603.00184v1[cs.CV] 26 Feb 2026

### v1

@@ Abstract
I Introduction
@@
II Related Work
II-A End-to-End Segmentation Networks
@@ II-B Instance Segmentation
II-C Foundation Models for Segmentation
@@
II-D Open Vocabulary Detection
@@ II-E YOLO Series
II-F Bird Specific Prior Work
@@
III Background: Foundation Models
III-A SAM 2.1
@@ III-C YOLOv11
IV Methodology
IV-A System Architecture
@@
IV-B Pipeline A: Zero-Shot (Grounding DINO 1.5 + SAM 2.1)
@@
IV-C Pipeline B: Supervised (YOLOv11 + SAM 2.1)
@@ IV-F Evaluation Metrics
V Experiments and Results
V-A Comparison with State-of-the-Art
Table I compares both our pipelines against six prior methods on the same CUB-200-2011 test split. The supervised YOLOv11 + SAM 2.1 pipeline outperforms all baselines on every metric, while the zero-shot pipeline (requiring no bird specific training) still exceeds SegFormer B2 on Dice and F1.
TABLE I: Comparison on CUB-200-2011 Test Set. $\dagger$ = zero-shot (no bird training). $\ddagger$ = detector-only training.
@@ V-A Comparison with State-of-the-Art
V-B Ablation: Effect of Detector Quality on SAM 2.1
Table II isolates the contribution of each detection stage while holding SAM 2.1 fixed. This demonstrates that SAM 2.1’s output quality is fundamentally gated by the quality of its bounding box prompts. With oracle ground truth boxes the pipeline achieves IoU 0.934 an upper bound showing further gains are possible with better detectors.
TABLE II: Ablation: Detection Stage vs. SAM 2.1 Output (SAM 2.1 weights fixed throughout).
@@
V-C Speed Analysis
Table III reports inference latency on a single NVIDIA A100 GPU. The supervised pipeline runs at 14 FPS sufficient for most ecological monitoring applications. MobileSAM can substitute SAM 2.1 for speed critical deployments at a modest 3 pp IoU cost.
TABLE III: Inference Speed on Single NVIDIA A100 GPU.
@@ V-C Speed Analysis
V-D Zero-Shot Per-Category Analysis
The zero-shot pipeline achieves IoU >> 0.85 on 143 of the 200 CUB species. Performance is weakest on small birds (hummingbirds, IoU $\approx$ 0.72) and heavily occluded species in dense foliage (IoU $\approx$ 0.68). For large, clearly visible birds (pelicans, albatrosses), IoU exceeds 0.91—comparable to the supervised pipeline.
VI Discussion
VI-A The Prompting Paradigm vs. End-to-End Training
@@ VI Discussion
VI-B Why SAM 2.1 Outperforms End-to-End Decoders
SAM 2.1’s mask decoder was trained on 1.1 billion masks—orders of magnitude more mask level supervision than any domain specific model. End-to-end trained models (SegFormer, U-Net) must learn both object localisation and precise boundary segmentation from limited labelled data. SAM 2.1 specialises in boundary precision; the detector handles localisation. Separation of concerns yields better outcomes for both subtasks.
VI-C Limitations

### v2

# Zero-Shot and Supervised Bird Image Segmentation Using Foundation Models: A Dual-Pipeline Approach with Grounding DINO 1.5, YOLOv11, and SAM 2.1

@@ Abstract
Index Terms—SAM 2.1, Segment Anything Model, Grounding DINO 1.5, YOLOv11, foundation models, zero-shot segmentation, bird segmentation, CUB-200-2011, bounding-box prompting.
I.  Introduction
@@
II.  Related Work
A. End-to-End Segmentation Networks
@@ B. Instance Segmentation
C. Foundation Models for Segmentation
@@
D. Open Vocabulary Detection
@@ E. YOLO Series
F. Bird Specific Prior Work
@@
III.  Background: Foundation Models
A. SAM 2.1
@@ C. YOLOv11
IV.  Methodology
A. System Architecture
@@
B. Pipeline A: Zero-Shot (Grounding DINO 1.5 + SAM 2.1)
@@
C. Pipeline B: Supervised (YOLOv11 + SAM 2.1)
@@ F. Evaluation Metrics
V.  Experiments and Results
A. Comparison with State-of-the-Art
Table 1 compares both our pipelines against six prior methods on the same CUB-200-2011 test split. The supervised YOLOv11 + SAM 2.1 pipeline outperforms all baselines on every metric, while the zero-shot pipeline (requiring no bird specific training) still exceeds SegFormer B2 on Dice and F1.
Table 1: Comparison on CUB-200-2011 Test Set. $\dagger$ = zero-shot (no bird training). $\ddagger$ = detector-only training.
@@ A. Comparison with State-of-the-Art
B. Ablation: Effect of Detector Quality on SAM 2.1
Table 2 isolates the contribution of each detection stage while holding SAM 2.1 fixed. This demonstrates that SAM 2.1’s output quality is fundamentally gated by the quality of its bounding box prompts. With oracle ground truth boxes the pipeline achieves IoU 0.934 an upper bound showing further gains are possible with better detectors.
Table 2: Ablation: Detection Stage vs. SAM 2.1 Output (SAM 2.1 weights fixed throughout).
@@
C. Speed Analysis
Table 3 reports inference latency on a single NVIDIA A100 GPU. The supervised pipeline runs at 14 FPS sufficient for most ecological monitoring applications. MobileSAM can substitute SAM 2.1 for speed critical deployments at a modest 3 pp IoU cost.
Table 3: Inference Speed on Single NVIDIA A100 GPU.
@@ C. Speed Analysis
D. Zero-Shot Per-Category Analysis
The zero-shot pipeline achieves IoU >> 0.85 on 143 of the 200 CUB species. Performance is weakest on small birds (hummingbirds, IoU $\approx$ 0.72) and heavily occluded species in dense foliage (IoU $\approx$ 0.68). For large, clearly visible birds (pelicans, albatrosses), IoU exceeds 0.91 comparable to the supervised pipeline.
VI.  Discussion
A. The Prompting Paradigm vs. End-to-End Training
@@ VI.  Discussion
B. Why SAM 2.1 Outperforms End-to-End Decoders
SAM 2.1’s mask decoder was trained on 1.1 billion masks orders of magnitude more mask level supervision than any domain specific model. End-to-end trained models (SegFormer, U-Net) must learn both object localisation and precise boundary segmentation from limited labelled data. SAM 2.1 specialises in boundary precision; the detector handles localisation. Separation of concerns yields better outcomes for both subtasks.
C. Limitations